# Thermodynamics of the dipole-octupole pyrochlore magnet CeHfO in applied magnetic fields

### v1

@@ IV Theoretical analysis for [110] magnetic fields
We now compare these results with the theoretical predictions based on our previously determined Hamiltonian parameters [57] reproduced in Eq. (2). These parameters were obtained from fitting our quantum many-body calculations to thermodynamic measurements in [111] magnetic fields and rescaled Landau-Lifshitz spin/molecular dynamics [70] to neutron scattering measurements in zero field. Since the data for [110] fields was not used for the fitting, a comparison between experiment and theory serves as a non-trivial check of our microscopic description.
@@
Theoretical results are obtained for both 16-site and 32-site pyrochlore clusters (with periodic boundary conditions) using the finite temperature Lanczos method (FTLM) [71] which has previously been applied for related systems [72, 56] (Further details of the method can be found in Ref. [71, 72]). While both cluster sizes are small, we see that their high temperature features coincide, a result expected when the correlation lengths are small (see red and blue curves in Fig. 3). This good agreement also extends to low temperatures, giving us confidence in the numerical results.
@@
Further, we note a very good agreement with the high temperature portion of the experimental data, reaffirming faith in the microscopic description and the parameter values. The high temperature bump is especially well reproduced suggesting that our value is well constrained. At lower temperature, however, there are quantitative differences with respect to the experiment. Calculations indeed predict two maxima in the specific heat (see Fig. 3 and dashed red curves in Figure 5), but the field evolution of the low temperature bump is not correctly captured.
@@
We also compare experimental magnetization curves with calculations based on parameters in Eq. 2 (Fig. 3(d)). Our theoretical curves (computed for the 16 sites pyrochlore cluster) agree well with the experiment for low fields at all temperatures, and for all fields at higher temperatures. At high fields and especially low temperatures, prominent deviations are observed, notably, the lack of saturation of the experimental magnetization. This regime is expected to be governed by the physics of a single tetrahedron (thereby ruling out significant finite size effects), with interactions playing only a diminished and subdominant role, thus it becomes imperative to scrutinize the origin of this discrepancy.
@@
Going back to specific heat under magnetic field along the [110] direction, the location of the (high temperature) Schottky bump originates from the spins that align themselves to minimize their Zeeman energy (ignoring the role of interactions). As the magnetic field strength is increased, the magnetic energy contribution from the chains increases and hence a higher temperature is required to completely disorder these spins, which is associated with the destruction of short range order. However, this does not address the underlying reason for the existence of the low temperature peak and its movement with increasing magnetic field. We investigate three scenarios, including one based on existing literature and our conclusions are the following:
(1) We consider the spins which, to linear order in the field strength, do not couple to the external magnetic field applied exactly along the direction. The spins are indirectly influenced by the field through their interaction with the chains, which can lead to their ordering through perturbative effects. However, this effect predicts a decrease in the location of the low temperature bump with increasing the field strength, a trend exactly opposite to that observed in experiment (setup B), ruling it out as an explanation of the data.
(2) We estimate the higher-order couplings between the octupolar spin components coupled to the magnetic field via a term . While this term can in principle yield a low temperature peak, the required is orders of magnitude larger than what we estimate from ab-initio CEF calculations. This mechanism is also ruled out, at least as a dominant mechanism.
(3) Finally, we consider the possibility that the external magnetic field is not exactly along the direction, so it can couple to the spins directly. We found this assumption explains the magnetic field dependence of the low temperature peak location, even for modestly small misalignment angles. Incorporation of small misalignments also largely resolves the discrepancy in the high-field magnetization.
@@ IV.1 First scenario: Indirect coupling of spins to a [110] magnetic field mediated by the spins
The spins couple to one another by field-dependent effective interactions, this is mediated via their interaction with the chains. This was placed in a theoretical framework by Placke, Benton and Moessner (PMB) [68] using a real space perturbation theory approach and order by disorder arguments [73]. While PMB extensively studied the classical phase diagram of dipole-octupole pyrochlores in a [110] magnetic field for a wide range of parameter sets, we will focus on the case where is dominant, since it is directly pertinent to the study of CeZrO and CeHfO. It is worth noting that the argumentation would essentially remain similar if is dominant.
The subextensive degeneracy is split by coupling between the chains, each representing an Ising spin, which themselves form an anisotropic triangular lattice. This coupling is a result of finite but small while keeping the [110] field large. In this regime, corresponding to the so-called ‘chain Y’ phase, PMB calculated the couplings between nearest neighbor Ising spins to be,
where decays as as .
@@
Importantly, this functional dependence suggests that the order-disorder transition of the chains should occur at lower temperatures upon increasing the field strength. This is opposite to the trend we observe in experiments, and thus we see no evidence of the PMB scenario in our current experiments.
We take this opportunity to clarify that the PMB procedure completely ignores quantum mechanical fluctuations within a chain. If these fluctuations are important, it is not apparent that it is justified to treat an entire chain as a single Ising degree of freedom. At large [110] fields the spins are essentially locked and can be treated classically. However, sites on a single chain are interacting quantum mechanically with their immediate nearest neighbors on the same chain, and this intrinsically selects a state on the scale of the magnetic coupling strengths (which are at the scale of 0.4 K or less), essentially independent of the magnetic field strength. We see evidence for our assertion in Fig. 4 where we plot the location of the low temperature feature as a function of field strength. We see only a small increase with increasing for low , and essentially independence at large .
@@
IV.2 Second scenario: Higher-order Zeeman coupling of pins to [110] field
@@
We now investigate the second scenario that enables a direct coupling of the spins with the applied [110] magnetic field by incorporating the next order Zeeman term in the effective Hamiltonian. This may seem odd at first, after all the bare ab-initio Hamiltonian has only a linear term. However, since we are dealing with a projected effective Hamiltonian restricted to the doublet subspace, we must admit the possibility of every symmetry allowed term including one that scales as .
The symmetry allowed third-order Zeeman term for the pseudo-spin 1/2 Hamiltonian is [63]
When the field is along [110] direction, , and (where is the field strength) and this makes the coefficient of in the above equation zero for sublattices 1 and 2. The contribution above (with only the term) has been previously considered in earlier work [63] but its quantitative impact on the low energy properties, to the best of our knowledge, has not been definitively established. We find here that this effect is too small to account for the observations of our experiments.
As a first exercise we attempted to fit the effective Hamiltonian parameters, along with the newly introduced term, to our specific heat data. While we found it to be somewhat large ( meV), and our fits could only reproduce the experimental features at low fields. In order to assess the possible importance of higher-order field terms and extract the directly, we performed diagonalization of the single ion ab-initio Hamiltonian which is a sum of the spin-orbit coupling term, the crystal field Hamiltonian and the Zeeman term in the space spanned by and manifold (i.e., Hilbert space of size 14). We especially ignore interactions between the ions. For a single ion, the form of the crystal field Hamiltonian is
Here , , , , and in meV. are the Stevens operators [74, 75]. These values of crystal field parameters were determined by fitting the CEF Hamiltonian parameters to inelastic neutron spectroscopy and magnetic susceptibility data measured on a powder sample of CeHfO as reported in [55]. The spin-orbit coupling part of the Hamiltonian is given by
Here, , which determines the energy scale of the spin-orbit interaction, was set at 80.3 meV. The bare Zeeman term is given by
The full Hamiltonian, for a given ion, is the sum of those terms
Next, a meaningful estimation of is performed by evaluating the energy splitting of the lowest lying doublet in the presence of [110] field using Eq. 10. For sublattices 1 and 2 we evaluate the energy difference between the first excited state and the ground state using Eq. 6 for each of them as a function of the magnetic field strength. By fitting to an appropriately defined cost function, this energy difference curve directly gives us , which we find to be meV (see Appendix A). Hence, the second scenario is not an appropriate mechanism to explain the observed two step release of entropy in CeHfO.
@@
The third scenario involves a direct Zeeman coupling of the chains to the magnetic field by assuming its misalignment away from the [110] direction. Since the preparation of specific heat measurements involves several steps subject to uncertainties (alignment from Laue diffraction images, sample cut, sample mounting on a given specific heat setup, and alignment of that setup relative to the superconducting solenoid), a slight misalignment of the applied magnetic field with respect to the required crystallographic orientation can hardly be avoided already at the start of measurements. Added to that, the torque induced by the application of a magnetic field on the sample may induce a deviation from the starting geometry. In particular, according to our experience, the measurement platform of the specific heat puck (commonly used for measurements in the PPMS) can tilt by several tens of degree at low temperature, without breaking the wires, when applying the field along the hard axis of a magnetic sample.
We now quantitatively determine the effect of misalignment on the specific heat data. To do so, we parametrize the global-basis components of a magnetic field tilted away from the [110] direction, as a function of its strength , polar angle , and azimuthal angle (see Fig. 1),
@@
Our attempts to explain the data with a single set of misalignment angles were not successful, so for each magnetic field strength we optimize the values of and independently to best fit the experimental [110] specific heat data. For each , the initial guess of and was chosen with and constrained beween 45 to 90 and 0 to 40 respectively. At each stage of the optimization (using the SLSQP algorithm in python) an FTLM calculation with seeds and (where is the number of independent Lanczos runs and is the dimension of the Krylov space) was performed. The cost function was designed to minimize the difference between the experimental and FTLM data.
@@
The results of our specific heat curves for the optimized misalignment angles and are shown in Fig. 6. Our optimized values of and suggest that the specific heat is more sensitive to values of rather than : the latter is found to be only a few degrees off from 90 degrees (fitting performed with being allowed to go beyond 90 also yielded an optimized value close to 90). We have thus shown the extent of variation of specific heat in Fig. 6 for the cases where was fixed and in the range . The optimal curves are shown in red and all the curves for different values of get the high temperature tail accurately. This is because at these temperatures the spins are disordered and tilting the field further does little to change the situation. Importantly, however, the different curves show variation in their lower temperature features based on which we infer that our has an error bar of a few degrees.
@@
Fields as large as 6 T are large enough for spin-spin interactions to be a sub-leading effect. In this regime, there are almost no finite size errors as the thermodynamics is largely governed by the response of four (almost) non-interacting sublattices. However, our optimized FTLM and experimental data do not perfectly match in this regime suggesting that effects beyond those incorporated in our model are required to better reproduce the experimental data.
@@
The qualitative trend emerging from our fits is that remains close to while increases with field strength. The increase of is steeper at low field and tends to saturate above 5 T, as seen in Fig. 6(f). To understand this behavior, we have calculated the torque expected for classical Ising moments on a tetrahedron as a function of field direction. The results are presented in Fig. 7 and show that the torque has a maximum for the [110] direction, around which it varies very differently depending on the direction of approach, or . Going away from the [110] direction in either or goes towards a minimum in torque corresponding to a lower Zeeman energy (Fig. 7, right panel). This energy landscape rationalizes the increase in deduced from the fits of the specific heat data. From an experimental viewpoint, both measurement setups for specific heat rely on thin wires to hold the sample, or sample platform, in the starting orientation, thus leaving room for movements under applied magnetic field. Although the fitted values from the high-field PPMS data appear relatively large, they are not inconsistent with our observations of widely tilted platform with other samples (see Appendix B).
@@
The emergent picture of the two bump feature in the specific heat in a misaligned sample is as follows. At low temperature, the spins choose an axis that is (almost) the local + or - direction. Because of this locking and because the spins now see a direct field and also interact with the spins, they too choose a unique ground state from an extensively degenerate manifold. Once the temperature is increased, the spins disorder first, yielding the first bump. Upon increasing the temperature further, the spins disorder, which corresponds to the second bump. These scales are distinct at low field strengths where the misalignment is small. At high field strengths the misalignment is larger, and the Zeeman energy scales for and spins become comparable. The double bump feature thus gradually fades away to yield a specific heat with a single bump as the field is increased, an effect that is indeed captured by our FTLM simulations.
@@
In Fig. 8 we quantitatively check our theoretical results for the two step release of entropy with respect to the experimental results. The experimental results (where the peak could be clearly discerned) suggest that both peak locations move linearly with field strength. This broadly supports the assertion that the sublattice also experiences a direct magnetic field in accordance with the misalignment theory. The location of the high-temperature bump calculated with FTLM is in excellent agreement with experiments, but the low temperature feature, while showing a generally correct qualitative trend, displays some quantitative discrepancy.
@@
Lastly, we also explore the effect of misalignment on the magnetization. It should be noted that the experimental magnetization has a gradual, but non zero, slope at large values of field strength, which is a regime primarily dominated by physics of a single tetrahedron. Somewhat surprisingly, our theoretical calculations failed to capture this slope. Further investigations were carried out to probe the role of impurities in the sample - we found that their effects are likely too small to explain the magnitude of the feature seen in the experiment. We refer the reader to Appendix C for further discussion on the role of impurities.
@@
Calculations from ab-initio calculations (using PyCEF and SPECTRE) working in a Hilbert space of 14 states, i.e. keeping all the and states, yielded a small but non-negligible van Vleck contribution to the susceptibility, . This non-zero contribution emerges from the presence of higher crystal field levels beyond the doublet, and results in a non-zero slope in the versus curve at large fields. (This latter feature does not exist in the pseudo spin-1/2 model with only a linear Zeeman term.) Despite its qualitative correctness, quantitatively the calculated is a factor of 5 smaller than what is needed to explain the slope seen in the magnetization experiment.

### v2

@@ IV Theoretical analysis for [110] magnetic fields
We now compare our experimental results with the theoretical predictions based on our previously determined Hamiltonian parameters [61] reproduced in Eq. (2). These parameters were obtained from fitting quantum many-body calculations to thermodynamic measurements in [111] magnetic fields and rescaled classical Landau-Lifshitz spin/molecular dynamics [72] to neutron scattering measurements in zero field. Since the data for [110] fields was not used for the fitting, a comparison between experiment and theory serves as a non-trivial check of the microscopic description.
@@
Theoretical results are obtained for both 16-site and 32-site pyrochlore clusters (with periodic boundary conditions) using the finite temperature Lanczos method (FTLM) [73] which has previously been applied for related systems [71, 56] (Further details of the method and its application to the pyrochlore geometry can be found in Ref. [73, 71]). While both cluster sizes are small, we see that their high temperature features coincide, a result expected when the correlation lengths are short (see red and blue curves in Fig. 3 (a-c)). This good agreement also extends to lower temperatures, giving us confidence in the numerical results. Our calculations also suggest that finite size effects, which are minimally present at higher temperatures, do become more prominent at lower temperatures, especially at lower fields, see for example the results for 0.5 T in Fig. 3(a). This is anticipated given the existence of the putative gapless spin liquid at zero field [61]. However, in practice even a modest field strength ( 1 T) appears to be dominated by the physics of a single tetrahedron, which is why we employ the 16-site cluster only for the rest of the paper.
@@
We also highlight the agreement of the computed specific heat with the high temperature portion of the experimental data, reaffirming faith in the theoretical microscopic description and the parameter values, specifically suggesting that the value is well-constrained. At lower temperature, however, there are quantitative differences with respect to the experiment. Calculations indeed predict two peaks in the specific heat (see Fig. 3 (a-c) and dashed red curves in Figure 4), but the field evolution of the low temperature peak is not correctly captured.
@@
We also compare experimental magnetization curves with calculations based on parameters in Eq. 2 (see Fig. 3(d)). Our theoretical curves (computed for the 16 site pyrochlore cluster) agree well with the experiment for low fields at all temperatures, and for all fields at higher temperatures. At high fields and especially low temperatures, prominent deviations are observed, notably, the lack of saturation of the experimental magnetization. As mentioned above, this regime is expected to be governed by the physics of a single tetrahedron, thereby ruling out significant finite size effects, with interactions playing only a diminished and subdominant role. Thus it becomes imperative to scrutinize the origin of this discrepancy.
@@
We revisit the temperature dependence of the specific heat in a [110] magnetic field. For large fields, the location of the (high temperature) Schottky bump originates from the spins that align themselves to minimize their Zeeman energy. As the magnetic field strength is increased, the magnetic energy contribution from the chains increases and hence a higher temperature is required to completely disorder these spins, which is associated with the destruction of short-range order. However, this does not address the underlying reason for the existence of the low temperature feature and its movement with increasing magnetic field. To address this we investigated three scenarios, including one based on existing literature, which we discuss next.
@@ IV.1 First scenario: Indirect coupling of spins to a [110] magnetic field mediated by the spins
The spins couple to one another by field-dependent effective interactions, this is mediated via their interaction with the chains. This was placed in a theoretical framework by Placke, Benton and Moessner (PMB) [53] using a real space perturbation theory approach and order by disorder arguments [74] treating each chain as an effective Ising degree of freedom. For dominant and small while keeping the [110] field large, PMB calculated the couplings between nearest neighbor Ising spins to be,
@@
where decays as as . This functional dependence suggests that the order-disorder transition of the chains should occur at lower temperatures upon increasing the field strength. This is opposite to the trend we observe in experiments, and thus we see no evidence of the PMB scenario in our current experiments. We discuss this further in Appendix A.1.
@@
IV.2 Second scenario: Higher-order Zeeman coupling of spins to [110] field
@@
We also explored the possibility of a direct coupling of the spins with the applied [110] magnetic field by incorporating the next order Zeeman term [52]. This is allowed by symmetry for the projected effective Hamiltonian, restricted to the doublet subspace, and the term is given by
When the field is along [110] direction, , and (where is the field strength) and this makes the coefficient of in the above equation zero for sublattices 1 and 2. In Appendix A.2 we discuss the details of our estimation of from ab-initio (CEF) calculations. We find it to be small meV-2, this smallness rules out this mechanism as a quantitative explanation of the two step release of entropy in Ce2Hf2O7.
@@
The third scenario involves a direct Zeeman coupling of the chains to the magnetic field by assuming its misalignment away from the [110] direction. Since the preparation of specific heat measurements involves several steps subject to uncertainties (alignment from Laue diffraction images, sample cut, sample mounting on a given specific heat setup, and alignment of that setup relative to the superconducting solenoid), a slight misalignment of the applied magnetic field with respect to the required crystallographic orientation can hardly be avoided already at the start of measurements. Added to that, the torque induced by the application of a magnetic field on the sample may induce a deviation from the starting geometry. In particular, according to our experience, the measurement platform of the specific heat puck (commonly used for measurements in the PPMS) can tilt by several tens of degrees at low temperature, without breaking the wires, when applying the field along the hard axis of a magnetic sample.
We now quantitatively determine the effect of misalignment on the specific heat data. To do so, we parameterize the global-basis components of a magnetic field tilted away from the [110] direction, as a function of its strength , polar angle , and azimuthal angle (see Fig. 1),
@@
Our attempts to explain the data with a single set of misalignment angles were not successful, so for each magnetic field strength we optimize the values of and independently to best fit the experimental [110] specific heat data. For each , the initial guess of and was chosen with and constrained between 45∘ to 90∘ and 0∘ to 40∘ respectively. At each stage of the optimization (using the SLSQP algorithm in python) an FTLM calculation with seeds and (where is the number of independent Lanczos runs and is the dimension of the Krylov space) was performed , which were found to be sufficient for attaining convergence within the temperature range considered. The cost function was designed to minimize the difference between the experimental and FTLM data.
@@
The results of our specific heat curves for the optimized misalignment angles and are shown in Fig. 5(a-e). Our optimized values of and suggest that the specific heat is more sensitive to values of rather than : the latter is found to be only a few degrees off from 90 degrees (the fits that allowed to exceed 90 degrees also found the optimized value to be close to 90 degrees). We have thus shown the extent of variation of specific heat in Fig. 5 for the cases where was fixed and in the range . The optimal curves are shown in red and all the curves for different values of accurately capture the high temperature tail. This is because at these temperatures the spins are disordered and tilting the field further does little to change the situation. Importantly, however, the different curves show variation in their lower temperature features based on which we infer that our has an error bar of a few degrees.
@@
Owing to the sub-meV interaction scales, fields strengths of even a few Tesla are large enough for spin-spin interactions to be a sub-leading effect. For example, at T there are almost no finite size errors as the thermodynamics is largely governed by the response of four (almost) non-interacting sublattices. However, our optimized FTLM and experimental data do not perfectly match in this regime suggesting that effects beyond those incorporated in our model are required to better reproduce the experimental data.
@@
The qualitative trend from our fits suggests remains close to while increases with field strength. The increase of is steeper at low field and tends to saturate above 5 T, as seen in Fig. 5(f). To understand this behavior, we have calculated the torque expected for classical Ising moments on a tetrahedron as a function of field direction. The results are presented in Fig. 6 and show that the torque has a maximum for the [110] direction, around which it varies very differently depending on the direction of approach, or . Going away from the [110] direction in either or goes towards a minimum in torque corresponding to a lower Zeeman energy (Fig. 6(b)). This energy landscape rationalizes the increase in deduced from the fits of the specific heat data. From an experimental viewpoint, both measurement setups for specific heat rely on thin wires to hold the sample, or sample platform, in the starting orientation, thus leaving room for movements under applied magnetic field. Although the fitted values from the high-field PPMS data appear relatively large, they are broadly consistent with our observations of widely tilted platform with other samples (see Appendix B).
@@
The emergent picture of the two peak feature in the specific heat in a misaligned sample is as follows. At low temperature and high fields, the spins choose an axis that is (almost) the local + or - direction. Because of this locking and because the spins now see a small but direct field and also interact with the spins, they too choose a unique ground state from an extensively degenerate manifold. Once the temperature is increased, the spins disorder first, yielding the first peak. Upon increasing the temperature further, the spins disorder, which corresponds to the second bump. These scales are distinct at low field strengths where the misalignment is small. At high field strengths the misalignment is larger, and the Zeeman energy scales for and spins become comparable. The double peak feature thus gradually fades away to yield a specific heat with a single bump as the field is increased, an effect that is indeed captured by our FTLM simulations.
@@
In Fig. 7 we quantitatively check our theoretical results for the two step release of entropy with respect to the experimental results. The experimental results (where the peak could be clearly discerned) suggest that both peak locations move linearly with field strength. This broadly supports the assertion that the sublattice also experiences a direct magnetic field in accordance with the misalignment theory. The location of the high-temperature bump calculated with FTLM is in excellent agreement with experiments, but the low temperature feature, while showing a generally correct qualitative trend, displays some quantitative discrepancy. We also note in Fig. 7 that the slope of the low-temperature peak changes abruptly past 2 T, which is due to data from two different setups (setup A at low fields and setup B at higher fields), but no such difference is present in the high-temperature bump. We attribute this to the location of the high temperature bump being less susceptible to misalignment (which we expect to be different in the two setups), as it arises from melting of the spins which directly couple to the [110] magnetic field.
@@
Lastly, we also explore the effect of misalignment on the magnetization. Note that the experimental magnetization has a gradual, but non zero, slope at large values of field strength, which is a regime primarily dominated by physics of a single tetrahedron. Further investigations were carried out to probe the role of impurities in the sample - we found that their effects are likely too small to explain the magnitude of the feature seen in the experiment. (We refer the reader to Appendix C for discussion on the role of impurities.)
@@
Calculations from ab-initio calculations (using PyCEF [75] and SPECTRE [76]) working in a Hilbert space of 14 states, i.e. keeping all the and states, yielded a small but non-negligible van Vleck contribution to the susceptibility, . This non-zero contribution emerges from the presence of higher crystal field levels beyond the doublet, and results in a non-zero slope in the versus curve at large fields. (This latter feature does not exist in the pseudo spin-1/2 model with only a linear Zeeman term.) Despite its qualitative correctness, quantitatively the calculated is a factor of 5 smaller than what is needed to explain the slope seen in the magnetization experiment.